OpenStudy (anonymous):

I guess if you guys get tired of me asking you'll quit answering, LOL!! So this one is too weird for me: 6-2(3x-3)=0. I get X=0. Can someone tell me if I am right?

OpenStudy (anonymous):

nono well keep answering :)!

OpenStudy (anonymous):

;) thanks

OpenStudy (anonymous):

x = 0 you are correct

OpenStudy (anonymous):

REALLY!! Wow that's great, thank you much for your help:)

OpenStudy (anonymous):

you can also plug x-0 back into the equation and after some algebra you will see 0=0 so therfore an x value of 0 works

OpenStudy (anonymous):

Ok, I don't quite get that but I'll learn:)

OpenStudy (amistre64):

6+6 not= 0 tho

OpenStudy (anonymous):

Ok, thank you for that Amistre:)

OpenStudy (amistre64):

youre welcome, but if you wnt to step thru it id be willing to help out 6-2(3x-3)=0 ; subtract 6 from both sides -6 -6 ---- ---- -2(3x-3) = -6 ; divide off that -2 /-2 /-2 ----- ---- 3x-3 = 3 ; add 3 to each side +3 +3 ----- --- 3x = 6

OpenStudy (anonymous):

Oh, you mean the answer isn't x=0?

OpenStudy (amistre64):

the answer is not x=0

OpenStudy (anonymous):

oh wow!!!

OpenStudy (anonymous):

I know the procedure you used and I know it's right. I usually ask folks to show me how they got the answer Thank you much!!!

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

so is 3x =3 the simplified answer, as far as we can go with it?